# arXiv:2603.01298v1[q-fin.PM] 01 Mar 2026

### v1

@@ Acknowledgments
Appendix A Evaluation
Suppose we produced an index that had a volatility exactly equal to $\sigma^{\text{tar}}$. Any empirical estimate of the volatility $\sigma^{\text{eval}}_{k}$ from realized returns will almost surely not equal $\sigma^{\text{tar}}$. It is thus important to understand how closely we can expect our volatility estimates to match the target volatility.
For this purpose, we suggest the following benchmarking methodology:
@@ Appendix A Evaluation
    Independently sample $N$ (for some large integer $N$) synthetic return values $r_{1}^{\text{synth}},\ldots,r_{N}^{\text{synth}}$ from a normal distribution with mean 0 and standard deviation $\sigma^{\text{tar}}$.
@@
    Compute volatility estimates $\sigma^{\text{synth}}_{k}$ from the synthetic returns $r_{1}^{\text{synth}},\ldots,r_{k}^{\text{synth}}$ using the same method we used to compute $\sigma^{\text{eval}}_{k}$ from the index returns $r_{1}^{\text{ind}},\ldots,r_{k}^{\text{ind}}$.
@@
    Discard the first $n$ values of $\sigma^{\text{synth}}_{k}$ (where $n$ is a burn-in period for the volatility estimate).
  4. 4.
@@
    Compute statistics of interest from the remaining values of $\sigma^{\text{synth}}_{n+1},\ldots,\sigma^{\text{synth}}_{N}$.
Specifically, because we do not have the ability to observe the true volatility of our index, this methodology allows us to quantify the uncertainty of volatility estimates and understand the fluctuations we can expect to see if the index is tracking the target volatility.
A.1 Example
Suppose we estimate the volatility of our index via a EWMA with halflife $h$. We define $\alpha=\exp(-\log(2)/h)$ to be the corresponding decay factor of the EWMA. If we had infinitely many synthetic returns values from a normal distribution with mean 0 and standard deviation $\sigma^{\text{tar}}$, then we can define the infinite-sample, halflife-$h$ EWMA volatility estimate
$\sigma^{\text{synth}}=\sqrt{(1-\alpha)\sum_{i=1}^{\infty}\alpha^{i-1}\left(r_{i}^{\text{synth}}\right)^{2}}.$
It is a result of probability that $\sigma^{\text{synth}}$ is approximately distributed according to a scaled $\chi$ distribution with $\frac{1-\alpha^{2}}{(1-\alpha)^{2}}$ degrees of freedom and a scale factor of $\sigma\frac{1-\alpha}{\sqrt{1-\alpha^{2}}}$. From this approximation, we can estimate the standard deviation of these of EWMA volatility estimates as a function of the halflife $h$. For $\sigma^{\text{tar}}=0.15/\sqrt{252}$, corresponding to an annualized volatility of $15\%$, we show the estimated standard deviation of $\sigma^{\text{synth}}$ as a function of $h$ in Figure 3.
@@ A.1 Example
Figure 3: Approximate standard deviation of the ideal EWMA volatility estimate $\sigma^{\text{synth}}$ as a function of halflife $h$ for $\sigma^{\text{tar}}=0.15/\sqrt{252}$.
Returning to the sampling-based methodology we proposed, we can verify that for reasonably large $N$, the synthetic volatility estimates $\sigma^{\text{synth}}_{n+1},\ldots,\sigma^{\text{synth}}_{N}$ will be distributed according to the distribution of $\sigma^{\text{synth}}$. For $\sigma^{\text{tar}}=0.15/\sqrt{252}$ and halflives $h=21,126$, we demonstrate this in Figure 4.
@@
Figure 4: Distribution of EWMA volatility estimates $\sigma^{\text{synth}}$ from Monte Carlo simulation with $N=10000$ samples and the first $n=252$ discarded. The dashed lines show the approximate distribution of $\sigma^{\text{synth}}$ for halflives $h=21,126$.
@@
As evidenced by Figures 3 and 4, EWMA volatility estimates of $\sigma^{\text{tar}}$ with longer halflives have significantly less variance. As such, we recommend using halflives of at least 63 when evaluating the performance of our index construction methods.

### v2

@@ Acknowledgments
Appendix A Uncertainty Quantification for Volatility Estimates
Suppose we have constructed an index with returns $r_{1}^{\text{ind}},r_{2}^{\text{ind}},\ldots$ and we estimate its volatility using some method (e.g., the EWMA defined in (2)) to obtain estimates $\sigma^{\text{eval}}_{1},\sigma^{\text{eval}}_{2},\ldots$. Even if the index had a volatility exactly equal to the target $\sigma^{\text{tar}}$, any finite-sample estimate $\sigma^{\text{eval}}_{k}$ will almost surely differ from $\sigma^{\text{tar}}$ due to sampling variability. To evaluate whether our index is successfully tracking the target, we must understand how much deviation to expect from a perfectly tracking index.
We propose the following Monte Carlo methodology to quantify this uncertainty:
@@ Appendix A Uncertainty Quantification for Volatility Estimates
    Sample $N$ synthetic return values $r_{1}^{\text{synth}},\ldots,r_{N}^{\text{synth}}$ independently from a normal distribution with mean 0 and standard deviation $\sigma^{\text{tar}}$.
@@
    Compute volatility estimates $\sigma^{\text{synth}}_{1},\ldots,\sigma^{\text{synth}}_{N}$ from the synthetic returns using the same estimation method used to compute $\sigma^{\text{eval}}_{k}$ from the index returns.
@@
    Discard the first $n$ values (a burn-in period) and compute statistics of interest from $\sigma^{\text{synth}}_{n+1},\ldots,\sigma^{\text{synth}}_{N}$.
Because the synthetic returns have true volatility $\sigma^{\text{tar}}$ by construction, the distribution of $\sigma^{\text{synth}}_{k}$ characterizes the sampling variability we would observe if our index were tracking the target perfectly. This distribution provides the confidence bands shown in Figure 2.
Appendix B Closed-Form Approximation for EWMA Estimators
While the Monte Carlo approach above is general, we can derive a closed-form approximation when using EWMA-based volatility estimates. We first consider the simpler case of a simple moving average (SMA) estimator, then extend to EWMA.
For an SMA estimator with window size $m$, the squared volatility estimate from i.i.d. normal returns with true standard deviation $\sigma$ is
$\left(\sigma^{\text{SMA}}\right)^{2}=\frac{1}{m}\sum_{i=1}^{m}\left(r_{i}\right)^{2}.$
Since each $r_{i}\sim\mathcal{N}(0,\sigma^{2})$, we have $(r_{i}/\sigma)^{2}\sim\chi^{2}_{1}$, and the sum of $m$ independent $\chi^{2}_{1}$ variables follows a $\chi^{2}_{m}$ distribution. Thus $(\sigma^{\text{SMA}})^{2}\sim(\sigma^{2}/m)\chi^{2}_{m}$, or equivalently, $\sigma^{\text{SMA}}$ follows a scaled chi distribution with $m$ degrees of freedom and scale $\sigma/\sqrt{m}$.
For an EWMA estimator with decay factor $\beta=\exp(-\log(2)/h)$ corresponding to halflife $h$, the squared volatility estimate is a weighted sum of squared returns. In the infinite-sample limit,
$\left(\sigma^{\text{EWMA}}\right)^{2}=(1-\beta)\sum_{i=1}^{\infty}\beta^{i-1}r_{i}^{2}.$
This is no longer an exact chi-squared, but we can approximate its distribution by moment matching. Using $\mathbb{E}[r_{i}^{2}]=\sigma^{2}$ and $\text{Var}[r_{i}^{2}]=2\sigma^{4}$ (since $r_{i}^{2}/\sigma^{2}\sim\chi^{2}_{1}$), we obtain
$\mathbb{E}[(\sigma^{\text{EWMA}})^{2}]$ $=\sigma^{2},$
$\text{Var}[(\sigma^{\text{EWMA}})^{2}]$ $=(1-\beta)^{2}\frac{2\sigma^{4}}{1-\beta^{2}}=\frac{2\sigma^{4}(1-\beta)}{1+\beta}.$
A scaled chi-squared $(\sigma^{2}/\nu)\chi^{2}_{\nu}$ has mean $\sigma^{2}$ and variance $2\sigma^{4}/\nu$. Matching variance gives the effective degrees of freedom $\nu=(1+\beta)/(1-\beta)$. Thus $\sigma^{\text{EWMA}}$ is approximately distributed as a scaled chi distribution with $(1+\beta)/(1-\beta)$ degrees of freedom and scale $\sigma\sqrt{(1-\beta)/(1+\beta)}$.
@@ Appendix B Closed-Form Approximation for EWMA Estimators
Figure 6 shows the approximate standard deviation of $\sigma^{\text{EWMA}}$ as a function of halflife $h$ for $\sigma^{\text{tar}}=0.15/\sqrt{252}$ (annualized volatility of $15\%$).
Figure 6: Approximate standard deviation of the EWMA volatility estimate as a function of halflife $h$ for $\sigma^{\text{tar}}=0.15/\sqrt{252}$.
Figure 7 validates this approximation by comparing the closed-form chi distribution to the empirical distribution from Monte Carlo simulation with $N=10000$ samples.
@@
Figure 7: Histogram of EWMA volatility estimates from Monte Carlo simulation ($N=10000$, burn-in $n=252$) compared to the chi distribution approximation (dashed lines) for halflives $h=21$ and $h=126$.
@@
Figures 6 and 7 show that EWMA volatility estimates with longer halflives have significantly less variance. As such, we recommend using halflives of at least 63 when evaluating the performance of our index construction methods.